# Do Larger Language Models Imply Better Generalization? A Pretraining Scaling Law for Implicit Reasoning

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks requiring complex reasoning. However, the effects of scaling on their reasoning abilities remain insufficiently understood. In this paper, we introduce a synthetic multihop reasoning environment designed to closely replicate the structure and distribution of real-world large-scale knowledge graphs. Our reasoning task involves completing missing edges in the graph, which requires advanced multi-hop reasoning and mimics real-world reasoning scenarios. To evaluate this, we pretrain language models (LMs) from scratch solely on triples from the incomplete graph and assess their ability to infer the missing edges. Interestingly, we observe that overparameterization can impair reasoning performance due to excessive memorization. We investigate different factors that affect this U-shaped loss curve, including graph structure, model size, and training steps. To predict the optimal model size for a specific knowledge graph, we find an empirical scaling that linearly maps the knowledge graph search entropy to the optimal model size. This work provides new insights into the relationship between scaling and reasoning in LLMs, shedding light on possible ways to optimize their performance for reasoning tasks.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks. Recently, the reasoning capacity of LLMs has drawn a lot of attention as it is highly correlated with LLMs’ performance on many complex real-world tasks (Wei et al., 2022a; Guo et al., 2025). While the reasoning capability is usually enhanced during the post-training stage, it is reasonable to assume that LLMs have already acquired the capability during the pretraining stage, as the post-training is on a significantly smaller scale than pretraining. Some recent work has investigated the possible mechanism of LLMs learning to reason through next-token prediction pretraining (Zhu et al., 2024; Wang et al., 2024a; b). However, the impact of pretraining scaling on LLMs’ reasoning ability remains insufficiently understood.
@@
The general scaling behavior of language models has been investigated, including the well-known exponential scaling laws for testing loss and compute proposed by Kaplan et al. (2020) and the training compute-optimal scaling studied by Hoffmann et al. (2022a). Recent work has also examined the scaling of specific capabilities like machine translation (Ghorbani et al., 2022) and knowledge capacity/memorization (Allen-Zhu & Li, 2025; Lu et al., 2024). According to these existing scaling laws, it is in general believed that larger models imply better testing loss or task performance when trained on more data.
@@
In this paper, however, we find that the reasoning ability of language models under pretraining scaling can behave differently from normal power-law scaling, in a simplified pretraining environment. More specifically, when given enough compute, we find that the testing loss scaling curve is U-shape and there exists an optimal model size that produces the best reasoning performance/testing loss. This implies that overparameterization might hurt reasoning capability during pre-training. We first observe this phenomenon with real-world knowledge graph data, and then systematically study it through synthetically generated data.
We choose to mimic the real-world knowledge structure and distribution with synthetic knowledge graphs (KGs). We define reasoning over world knowledge as completing missing edges in an incomplete knowledge graph, which requires multiple jumps on the knowledge graph according to some pre-defined rules that are latently encoded into the graph generation process. To analyze this, we pretrain LMs from scratch using only triples from the incomplete graph and evaluate their ability to infer missing connections.
@@
We investigate important factors that affect the U-shape scaling of reasoning loss versus language model size. Our important findings can be summarized as follows:
@@
  • The minimum reasoning loss/maximum reasoning accuracy that a language model can reach is capped by the training data, regardless of the training steps and model size.
  • The optimal model size for a training corpus is largely fixed regardless of the training steps when the number of training steps is large enough.
  • When the underlying knowledge graph is fixed, training on more data sampled from the graph increases the optimal model size and reasoning performance.
  • More complex knowledge graph implies a larger optimal model size.
As we observed that the optimal model size is likely solely determined by the training knowledge graph, we then aim to find an empirical scaling law that can predict the optimal model size with knowledge graph statistics. We then discover a linear relationship between the optimal model size and a newly proposed graph search entropy, which measures the entropy of performing random searches on a knowledge graph. Roughly, 124 additional parameters in the optimal model size are required per 1-bit entropy increase in searching a knowledge graph.
@@
Our work contributes to the broader understanding of LLM reasoning by shedding light on the intricate relationship between scaling and reasoning capability. Our proposed empirical reasoning scaling law provides possible practical insights for optimizing LLMs’ reasoning ability at pretraining time.
2 Preliminaries
@@
While the real-world LLMs are pretrained on large scale text corpus, this corpus can be viewed as encoding a wide range of world knowledge. The power of LLMs lies in the fact that they can not only memorize the world knowledge and extract the knowledge when queried, but also reason over the world knowledge and draw novel conclusions. In this paper, we propose to construct a simplified pretraining corpus directly from a knowledge graph. A knowledge graph is comprised of a set of triples, and we use each knowledge triple as a training example. We test the reasoning capability of a language model trained on such a corpus by testing its accuracy in completing triples that have never been seen in the knowledge graph but can be deduced through latent rules encoded in the graph structure. For example, if we know A is B’s father, and B is C’s father, then we can deduce that A is C’s grandfather.
@@
To eliminate confounding variables and information contained in the lexical form of the entity and relation names, we label each entity and relation with a random ID and tokenize the IDs by characters. We use the LlaMA (Touvron et al., 2023) model architecture to implement LMs of different sizes by adjusting the hidden dimensions and the number of layers. The specific parameter scheme can be found in the Appendix A.
To evaluate the language model’s capability of reasoning over the knowledge graph, we test the language models on a held-out set of triples that are not seen in the training time. Note that all entity and relation types should have been seen during training time and the language model is only tasked to connect missing edges. To eliminate the need to generate the correct form of relation and entity IDs, and to handle the case where multiple correct answers exist, we design the testing set to be 10-option multiple-choice questions: the language model is tasked to choose the correct tail entity given the head entity and the relation. We ensure that there is only one correct answer among the given 10 options. Suppose there are $M$ questions in the testing set. [^1] We fix $M=1000$ for all of our experiments. For a ground truth triple $(e^{h},r,e^{t})$, we design 9 distracting options $e^{(1)},e^{(2)},...,e^{(9)}$. Then we use the test accuracy $\text{Acc}(\theta,G)$ and testing loss $\ell(\theta,G)$ to evaluate the reasoning capability of a language model $\theta$ over the knowledge graph $G$:
@@
3 Real-world Experiments
@@
Figure 1: The multiple-choice accuracy/loss on unseen triples of different-sized language models trained on a real-word knowledge graph FB15K-237. The left panel (trained with 10k steps) shows that the testing accuracy decreases after a certain model size. The middle panel shows U-shape loss curves of language models trained with different number steps. The right panel shows Note that the model size on x-axis is in log scale.
@@
In our initial sets of experiments, we investigate the reasoning scaling effect using a real-world knowledge graph, FB15K-237 (Toutanova & Chen, 2015). FB15K-237 is sampled from FB15K (Bordes et al., 2013), which is a dataset adapted from the Freebase knowledge base (Bollacker et al., 2007), a web-scale knowledge base released by Google. FB15K-237 contains $N_{e}=14,505$ entities, $N_{r}=237$ relations, and $N=310,116$ knowledge triples.
@@
In Figure 1, we show different-sized language models trained on FB15K-237 with different numbers of steps. When trained with the same number of steps, we observe a slight reasoning performance drop when using larger models. So we then look at the testing loss on these datasets and observe a U-shape trend with respect to the model size. This observation contradicts the previous belief that larger models yield a smaller testing loss. The training loss decreases monotonically with respect to model size.
@@
This implies that a language model can overfit to the training data when it is overparameterized for the underlying reasoning structure. Such deviation from traditional scaling law has also been reported in broken neural scaling law (Caballero et al., 2023) which proposed a double-descent-like (Nakkiran et al., 2020) function form instead of a monotonic power-law form. There have also been observations of tasks with inverse scaling (Wei et al., 2023) for large language models.
In this paper, we mainly focus on the scaling of model size. Instead of only scaling the size of the training data, we explore different possible ways of generating the knowledge graph and studying the effect of overall graph complexity on the model reasoning performance. In the following sections, we will mostly focus on understanding the ”turning point” of the reasoning loss. More specifically, we want to understand what is the optimal model size that can obtain the smallest possible reasoning testing loss.
@@
As shown in Figure 1 and in Section 5, we find this optimal model size is largely stable when training the model for enough steps. Note that, at training time, we repeat the training triples for many epochs (e.g. 30 times for FB15K-237) to find the optimal model size. This graph epoch is different from the real-world cases where we repeat the whole pretraining corpus for certain epochs. Because we can view each triple in the graph as a piece of factual knowledge (e.g. Barack Obama’s wife is Michelle Obama), this knowledge is usually repeated many times in a pretraining text corpus, in many different forms. Therefore, although our models have seen the same triple many times during training, the same piece of factual knowledge could also have been repeated several times in one pass of a real-world pretraining corpus.
@@ 4 Synthetic Data Construction
To investigate how the underlying knowledge structure influences language models’ reasoning performance, we propose an algorithm to generate synthetic knowledge graphs that mimic real-world knowledge graphs. More specifically, we assume that the knowledge graph generation process is governed by a set of logical rules.
@@
where $(e_{i-1},r_{i},e_{i})\in\mathcal{G}$. We abbreviate such rule by $h(r)=[r_{1},r_{2},...,r_{n}]$. We randomly generate a set of logical rules $\mathcal{H}$ and ensuring there is no cycles in the set. To grow a graph that follows these rules, we enforce sparsity of the possible relation types connecting to and branching out each entity. More specifically, we define node types based on the possible relation types connecting to and branching out each entity, based on the generated rules, as illustrated in Figure 2. Such sparsity is also observed in real-world knowledge graphs.
Our random graph generation process is inspired by the preferential attachment process (Barabási & Albert, 1999), which is used for generating scale-free networks with a power-law distribution for the degrees of the nodes. Intuitively, preferential attachment implies a “the rich get richer” approach to edge placement in the graph. Each time a new node is added to the graph, there is a ‘preference’ to connect to the nodes that are already highly connected, with a probability proportional to the target node’s degree. Since we have observed the scale-free property in real-world knowledge graphs and the internet is known to be a scale-free network, we adopt a preferential attachment based graph generation process. To accommodate different relation types assigned to each edge, we maintain a degree distribution for each relationship and add new edges according to preferential attachment.
The code for our random graph generation algorithm is shown in the Appendix B. In summary, we first randomly generate a set of rules $\mathcal{H}$, with the number of rules $|\mathcal{H}|=N_{h}$ and the range of rule length $[L_{min},L_{max}]$ as hyperparameters. Then we generate all possible node types as illustrated in Figure 2, with the maximum number of relations per node $M_{r}$ as a hyperparameter. We generate a seed graph by instantiating each rule with a set of new entities. To this, we incrementally add one new entity until the number of entities reaches $N_{r}$, by first randomly assigning a node type to it, and then randomly sampling the $m$ relation types from the set of relations defined by the node type. We choose the target of these $m$ new edges by preferential attachment. After adding every $K$ entities, we search through the current graph to add any edges that can be inferred through the logic rules defined in $\mathcal{H}$. We call the triples that can be deduced through a logic rule by deductible triples, otherwise atomic triples.
@@ 5.1 Graph Generation Ablation
We study the effects of the following four hyperparameters of graph data generation: the number of triples $N$, the number of entities $N_{e}$, the number of relations $N_{r}$, and the number of rules $N_{h}$. We fix all training hyperparameters as specified in the Appendix A but study the effect of training steps, as according to our preliminary experiments it has the largest effect on the optimal model size. The detailed data generation configuration for each set of experiments can also be found in the Appendix A.
Stable optimal model size with respect to training steps. In Figure 3 (a), we show the effect of training language models on the same knowledge graph with different numbers of training steps. As mention in the last part of Section 2, the optimal model size becomes smaller when the number of training steps increases, and then becomes stable after 4k steps. Another observation is regardless of the number of training steps, the maximum accuracy or minimum loss is stable. While we have ensured that all testing triples can be deduced through the training triples, there seems to be a performance cap determined solely by the knowledge graph data, which is unaffected by model size. In all following experiments, we train all models for 10k steps.

### v3

# Do Larger Language Models Generalize Better? A Scaling Law for Implicit Reasoning at Pretraining Time

@@ Abstract
Reasoning is an integral part of many tasks performed by language models (LMs). However, the effects of scaling model sizes and data on reasoning abilities at pretraining time remain understudied. To rigorously investigate this problem, we pretrain LMs from scratch on a synthetic implicit multihop reasoning environment designed to closely replicate the structure and distribution of real-world large-scale knowledge graphs. We then assess the LMs’ ability to complete the missing edges in the graph, which requires multi-hop reasoning that can be viewed as a simplification of implicit reasoning during real-world pretraining. Interestingly, we observe that overparameterization can impair the implicit reasoning performance due to excessive memorization. We investigate different factors that affect the loss curve when scaling different components of the knowledge graph, model size, and training steps. To predict the optimal model size for a specific knowledge graph, we find an empirical scaling law that shows optimal-sized LMs can approximately reason over 0.008 bit information per parameter. This work shows counterintuitive effects of model size scaling and provides new insights into the relationship between scaling and reasoning in LLMs.  [^1] Code: https://github.com/WANGXinyiLinda/reasoning-scaling-law
@@ 1 Introduction
Language Models (LMs) have demonstrated remarkable capabilities across a wide range of tasks, with reasoning being a core component (Wei et al., 2022a; Guo et al., 2025). While reasoning is typically enhanced during the post-training stage by encouraging LMs to generate long chain-of-thoughts (CoTs) (Guo et al., 2025; Yang et al., 2025), it is reasonable to assume that they already acquire the foundations of such capability during pretraining, given that post-training operates at a significantly smaller scale. Several recent studies have explored the mechanisms by which LMs may acquire reasoning-related abilities through next-token prediction pretraining (Zhu et al., 2024; Wang et al., 2024a; b). However, the impact of scaling on LMs’ reasoning ability during pretraining remains poorly understood.
@@
The general scaling behavior of LMs at pretraining time has been extensively investigated, including the well-known exponential scaling laws for testing loss and compute proposed by Kaplan et al. (2020) and the training compute-optimal scaling studied by Hoffmann et al. (2022a). Recent work has also examined the scaling of specific capabilities like machine translation (Ghorbani et al., 2022) and knowledge capacity/memorization (Allen-Zhu & Li, 2025; Lu et al., 2024). According to these existing scaling laws, it is in general believed that larger models imply better testing loss or task performance.
In this paper, instead we find that the scaling of LMs’ reasoning capability at pretraining time differs from normal power-law scaling, in a simplified controlled pretraining environment. We use implicit reasoning to denote the reasoning behavior that naturally emerges during pretraining. i.e. the capability to draw new conclusions from existing knowledge without being explicitly trained to generate chain-of-thoughts (CoTs). More specifically, we define implicit reasoning over world knowledge as the task of completing missing edges in an incomplete knowledge graph, which requires multi-hop traversal according to predefined logic rules that are implicitly encoded in the graph generation process. To investigate this, we pretrain LMs from scratch using only triples from the incomplete graph and then evaluate their ability to infer the missing connections.
@@
With sufficient compute, we find that the curve of implicit reasoning loss versus model size follows a U-shape, revealing an optimal model size that yields the best reasoning performance. This suggests that overparameterization may impair the implicit reasoning capability instilled during pretraining. We first observe this phenomenon using data derived from real-world knowledge graphs, and then systematically study it with synthetically generated data.
@@
We investigate important factors that affect the U-shaped scaling of reasoning loss versus language model size. Our important findings can be summarized as follows:
@@
  • The minimum reasoning loss reachable by an LM is solely determined by the training data, regardless of training steps and model size.
  • The optimal model size is solely determined by knowledge graph complexity and data size regardless of training steps.
  • We show that an optimal-sized LM can approximately reason over 0.008 bit information per parameter.
As we observed that the optimal model size is likely solely determined by the training knowledge graph, we then aim to find an empirical scaling law that can predict the optimal model size from knowledge graph statistics. We identify a linear relationship between the optimal model size and our proposed graph search entropy, which quantifies the entropy of performing random searches on a knowledge graph. Under this framework, we find that each parameter in the optimal model size can reason over approximately 0.008 bits of information in a knowledge graph. In contrast, Allen-Zhu & Li (2025) show that a language model can memorize up to 2 bits of information per parameter—substantially more than its reasoning capacity. This gap arises both from the greater difficulty of reasoning compared to memorization and from the different methodologies used to compute these information quantities. A more detailed discussion is provided in Section 5.2.
@@
Our work contributes to the broader understanding of LLM reasoning by shedding light on the intricate relationship between scaling and implicit reasoning capability. Our proposed empirical reasoning scaling law provides possible practical insights for optimizing LLMs’ implicit reasoning ability at pretraining time.
2 Method
@@
While real-world LLMs are pretrained on large scale text corpora, this corpus can be viewed as encoding a wide range of world knowledge. The power of LLMs lies in the fact that they can not only memorize the world knowledge and extract the knowledge when queried, but also reason over the world knowledge and draw novel conclusions. In this paper, we propose constructing a simplified pretraining corpus from a knowledge graph. A knowledge graph is comprised of a set of (head entity, relation, tail entity) triples, and we use each knowledge triple as a training example. We test the reasoning capability of a language model trained on such a corpus by testing its accuracy in completing triples that have never been seen in the knowledge graph but can be deduced through latent rules encoded in the graph structure. For example, if we know A is B’s father, and B is C’s father, then we can deduce that A is C’s grandfather.
@@
To eliminate confounding variables and information contained in the lexical form of the entity and relation names, we label each entity and relation with a random ID and tokenize the IDs by characters. We use the Llama (Touvron et al., 2023) model architecture to implement LMs of different sizes by adjusting the hidden dimensions and the number of layers. The specific parameter scheme can be found in the Appendix B.
To evaluate the language model’s capability of reasoning over the knowledge graph, we test the LMs on a held-out set of triples that are not seen in the training time. Note that all entity and relation types should have been seen during training time and the language model is only tasked to connect missing edges. To eliminate the need to generate the correct form of relation and entity IDs, and to handle the case where multiple correct answers exist, we design the testing set to be 10-option multiple-choice questions: the language model is tasked to choose the correct tail entity given the head entity and the relation. We ensure that there is only one correct answer among the given 10 options. Suppose there are $M$ questions in the testing set. [^2] We fix $M=1000$ for all of our experiments. For a ground truth triple $(e^{h},r,e^{t})$, we design 9 distracting options $e^{(1)},e^{(2)},...,e^{(9)}$. Then we use the test accuracy $\text{Acc}(\theta,G)$ and testing loss $\ell(\theta,G)$ to evaluate the reasoning capability of a language model $\theta$ over the knowledge graph $G$:
@@
3 Initial Experiments with Real-World Knowledge Graph
@@
Figure 1: The multiple-choice accuracy/loss on unseen triples of different-sized LMs trained on a real-word knowledge graph FB15K-237. The first column shows that the testing accuracy decreases after a certain model size. The second column shows U-shape loss curves of LMs trained with different numbers of steps. The third column shows the training loss decreases steadily. These trends are stable across different ways of processing the knowledge triples, with the triple-only data shows the cleanest trend. Note that the model size on x-axis is in log scale.
@@
In our initial sets of experiments, we investigate the reasoning scaling effect using a real-world knowledge graph, FB15K-237 (Toutanova & Chen, 2015). FB15K-237 is sampled from FB15K (Bordes et al., 2013), which is a dataset adapted from the Freebase knowledge base (Bollacker et al., 2007), a web-scale knowledge base released by Google. FB15K-237 contains $N_{e}=14,505$ entities, $N_{r}=237$ relations, and $N=310,116$ knowledge triples. We process this dataset in three different ways: (a) translate each knowledge triple into a natural language sentence by prompting GPT4 and then tokenize the sentence with a pre-trained tokenizer, as shown in the first row of Figure 1; (b) translate each knowledge triple into a natural language sentence using pre-generated templates, as show in the second row of Figure 1; (c) translate each knowledge triple into text by assigning a random ID to each entity and relation and tokenize them by characters, as shown in the last row of Figure 1. An example can be found in Appendix A Figure 5.
@@
In Figure 1, we show different-sized LMs trained on FB15K-237 in all settings with different numbers of training steps. We observe a consistant reasoning performance drop when using larger models, across different ways of processing the knowledge triples, while the training loss decreases monotonically with respect to model size. This observation contradicts the previous belief that larger models always yield a smaller testing loss.
@@
This implies that a language model can overfit to the training data when it is overparameterized for the underlying reasoning structure. Such deviation from traditional scaling law has also been reported in broken neural scaling law (Caballero et al., 2023) which proposed a double-descent-like (Nakkiran et al., 2020) function form instead of a monotonic power-law form. There have also been observations of tasks with inverse scaling (Wei et al., 2023) for large LMs.
In this paper, we focus primarily on the scaling of model size and data complexity. Rather than merely increasing the size of the training data, we explore many different setting for generating synthetic knowledge graphs. This allows us to ablate individual components of the graph generation process and examine how overall graph complexity affects reasoning. In the synthetic experiments presented below, we use random IDs instead of natural language sentences to eliminate lexical and syntactic effects, yielding cleaner trends from which we can draw quantitative conclusions.
@@
In the following sections, we will mostly focus on understanding the ”turning point” of the reasoning loss. More specifically, we want to understand what is the optimal model size, that is the model size that can obtain the smallest possible reasoning testing loss. As shown in Figure 1 and in Figure 3 (a), we find this optimal model size is largely stable when training the model for enough steps. Note that, at training time, we repeat the training triples for many epochs (e.g. 30 times for FB15K-237) to find the optimal model size. This graph epoch is different from the real-world cases where we repeat the whole pretraining corpus for certain epochs. Because we can view each triple in the graph as a piece of factual knowledge (e.g. Barack Obama’s wife is Michelle Obama), this knowledge is usually repeated many times in a pretraining text corpus, in many different forms. Therefore, although our models have seen the same triple many times during training, the same piece of factual knowledge could also have been repeated several times in one pass of a real-world pretraining corpus.
@@ 4 Synthetic Data Construction
To investigate how the underlying knowledge structure influences LMs’ reasoning performance, we propose an algorithm to generate synthetic knowledge graphs that mimic real-world knowledge graphs. More specifically, we assume that the knowledge graph generation process is governed by a set of logical rules.
@@
where $(e_{i-1},r_{i},e_{i})\in\mathcal{G}$. We abbreviate such rule by $h(r)=[r_{1},r_{2},...,r_{n}]$. We randomly generate a set of logical rules $\mathcal{H}$ and ensure there are no cycles in the set. To grow a graph that follows these rules, we enforce sparsity of the possible relation types connecting to and branching out each entity. More specifically, we define node types based on the possible relation types connecting to and branching out each entity, based on the generated rules, as illustrated in Figure 2. Such sparsity is also observed in real-world knowledge graphs.
Our random graph generation process is inspired by the preferential attachment process (Barabási & Albert, 1999), which is used for generating scale-free networks with a power-law distribution for the degrees of the nodes. Intuitively, preferential attachment implies a “the rich get richer” approach to edge placement in the graph. Each time a new node is added to the graph, there is a ‘preference’ to connect to the nodes that are already highly connected, with a probability proportional to the target node’s degree. Since we have observed the scale-free property in real-world knowledge graphs and the internet is known to be a scale-free network (Albert et al., 1999), we adopt a preferential attachment based graph generation process. To accommodate different relation types assigned to each edge, we maintain a degree distribution for each relationship and add new edges according to preferential attachment. A comparison of the node degree distribution between synthetic graph and real-world graph can be found in Appendix C Figure 6.
The code for our random graph generation algorithm is shown in the Appendix D. In summary, we first randomly generate a set of rules $\mathcal{H}$, with the number of rules $|\mathcal{H}|=N_{h}$ and the range of rule length $L_{min},L_{max}]$ as hyperparameters. Then we generate all possible node types as illustrated in Figure 2, with the maximum number of relations per node $M_{r}$ as a hyperparameter. We generate a seed graph by instantiating each rule with a set of new entities. To this, we incrementally add one new entity until the number of entities reaches $N_{r}$, by first randomly assigning a node type to it, and then randomly sampling the $m$ relation types from the set of relations defined by the node type. We choose the target of these $m$ new edges by preferential attachment. After adding every $K$ entities, we search through the current graph to add any edges that can be inferred through the logic rules defined in $\mathcal{H}$. We call the triples that can be deduced through a logic rule by deducible triples, otherwise atomic triples.
@@ 5.1 Graph Generation Ablation
We study the effects of the following four hyperparameters of graph data generation: the number of triples $N$, the number of entities $N_{e}$, the number of relations $N_{r}$, and the number of rules $N_{h}$. We fix all training hyperparameters as specified in the Appendix B. In all experiments except Figure 3 (a), we train all models for 10k steps. The detailed data generation configuration for each set of experiments can also be found in the Appendix B.
Stable optimal model size with respect to training steps. In Figure 3 (a), we show the effect of training language models on the same knowledge graph with different numbers of training steps. As mentioned in Section 3, the optimal model size becomes smaller when the number of training steps increases, and then becomes stable after 4k steps. Another observation is regardless of the number of training steps, the maximum accuracy or minimum loss is stable. While we have ensured that all testing triples can be deduced through the training triples, there seems to be a performance cap determined solely by the knowledge graph data, which is unaffected by model size.